Use the following financial statement information for Stowe Enterprises to solve Prob- lems 2A.4 through 2A.7: STOWE ENTERPRISES Statement of Comprehensive Income $4,500 2,400 500 Earnings before interest and taxes $1,600 480 $1,120 380 S 740 $ 186 554 2015 Sales Cost of goods sold Depreciation Interest paid Taxable income Taxes Net income Addition to retained earnings Dividends
STOWE ENTER PRISES Abbreviated Statement of Financial Position, 2014-2015 2014 2015 Liabilities and 2014 2015 Assets shareholders equity Current liabilities Current assets Cash Accounts receivable Inventory Fixed assets Net plant and equipment $ 800 $ 1,800 Accounts payable 1,550 1,630 1,629 1,380 1,950 3,135 2,040 2,300 Other Notes payable quity 3,487 625 570 1,146 746 Long-term debt Shareholders equity 4,620 5,130 common shares Accumulated retained earnings Total liabilities and 2,523 2.709 Total assets$10,505 $11270 shareholders equity $10,505 $11.270 Compute the following ratios for Stowe Enterprises for 2014 and 2015: Short-Term Asset Management Solvency Ratios Ratios Current ratio Total asset turnover Quick ratio Inventory turnover Cash ratio 2A4 Long-Term Solvency Ratios Debt ratio Debt-to-equity ratio Equity multiplier Interest coverage Cash coverage ratio Profitability Ratios Profit margin Return on assets Return on equity Receivables turnover 2A.5 2A.6 For how many days in 2015 could Stowe continue to operate if its production 2A.7 Stowe has 80 shares outstanding in 2015. The price per share is $45. What are 2A.8 Select an industry featured in recent financial news. Then obtain annual Prepare a statement of cash flows for Stowe. were suspended? the P/E ratio and the M/B ratio? reports on two companies in that industry and conduct a ratio analysis for the most recent two years. Make the relevant comparisons between the com panies and against industry norms. Baed on your ratio analysis, how do the two companies differ? Compare your comments against recent newspaper articles on the industry. (Note: This question is much easier to answer using ratio analysis software.)
Answer #1

As per rules I am answering the first 4 subparts of the question

A:Current Ratio = Current assets/ current liabilities

2014 = (800+1950+3135)/(1550+1629+746) = 1.5

2015 = (1800+2040+2300)/(1630+1380+625)= 1.69

B: Quick Ratio = (Cash + Receivables)/ Current liabilities

2014 = (800+1950)/(1550+1629+746) = 0.7

2015 = (1800+2040)/(1630+1380+625)= 1.06

C: Cash Ratio = Cash/ Current liabilities

2014 = 800/(1550+1629+746) = 0.2

2015 = 1800/(1630+1380+625)= 0.5

D: TOTAL asset turnover = Sales/ Average Total assets

2015 = 4500/ ((10505+11270)/2)

=41.33%